OpenStudy (zion):

between which numbers is log base8 256? 0 and 1 1and 2 2 and 3 64 and 256

OpenStudy (anonymous):

2 and 3

OpenStudy (zion):

how? solution pls.

OpenStudy (zarkon):

\[8^2=64\] \[8^3=512\]

OpenStudy (anonymous):

precisely. :D

OpenStudy (zion):

thanks!

OpenStudy (anonymous):

anyway, if youre wondering how.. see this \[2 ^{8} = 256\] and \[2 ^{3} = 8\] when youre doing log base 8, youre doing log base 2^3 ie dividing by 3log2. get it?

OpenStudy (anonymous):

i.e (8log2)/(3log2).. which gives 8/3. 3x2=6 &3x3=9. so 8/3 is more than 2 and less than 3.

OpenStudy (zion):

thanks alot.. u guys gave me hope that i will never be a noob in math,! :D